OpenStudy (anonymous):

Evaluate: 32 + (7 - 2) • 4 - 6 over 3 24 54 27 28 @ikram002p

OpenStudy (ikram002p):

\[32+(7-2)\times4 -\frac{ 6 }{ 3 } =32+(5\times4)-2=32+20-2=50\]